OpenStudy (anonymous):

find the perimeter of a rectangle that has a length of 3root2 + root 3 inches and a width of root 2 + root 3 inches

OpenStudy (anonymous):

\[8\sqrt{2}+4\sqrt{3}\] in inches

OpenStudy (anonymous):

thanks but can you please show your work? :)

OpenStudy (anonymous):

The perimeter of a rectangle is (2xL) + (2xW) \[3\sqrt{2}+\sqrt{3} = L\] \[\sqrt{2}+\sqrt{3} =W\] so 2 times L = \[6\sqrt{2}+2\sqrt{3}\] and 2 times W = \[2\sqrt{2}+2\sqrt{3}\] add them up = \[8\sqrt{2}+4\sqrt{3}\] correct me if I am wrong, hope it helps!